OpenStudy (phebe):

What is the solution of -3 = x + 5 A. -15 B. -8 C. 2 D. 8

OpenStudy (anonymous):

B

OpenStudy (phebe):

thanx can u help me with somehing

OpenStudy (phebe):

else

OpenStudy (anonymous):

i can try

OpenStudy (phebe):

ok thanx

OpenStudy (anonymous):

what is it ?

OpenStudy (phebe):

Which equation is equivalent to 2(3x – 1) – 3(5x – 3) = 4 ? A. -3x – 5 = 4 B. -9x – 4 = 4 C. -9x – 11 = 4 D. -9x + 7 = 4

OpenStudy (anonymous):

I can not see your question properly. May be a font problem on my side . Do you have a pic of that ?

OpenStudy (phebe):

ok hold on k

OpenStudy (phebe):

Which equation is equalvalent to : 2(3x-1)-3(5x-3)=4

OpenStudy (phebe):

A- -3x – 5 = 4 B- -9x – 4 = 4 C- -9x – 11 = 4 D- -9x + 7 = 4

OpenStudy (phebe):

@gemini208012

OpenStudy (phebe):

nvm i got it thas ok

OpenStudy (anonymous):

it will be -9x + 7 = 4

OpenStudy (phebe):

yea thas what i got thanx